# Knowledge-Augmented Multimodal Clinical Rationale Generation for Disease Diagnosis with Small Language Models

### v3

@@ Abstract
Interpretation is critical for disease diagnosis, but existing models struggle to balance predictive accuracy with human-understandable rationales. While large language models (LLMs) offer strong reasoning abilities, their clinical use is limited by high computational costs and restricted multimodal reasoning ability. Small language models (SLMs) are efficient but lack advanced reasoning for integrating multimodal medical data. In addition, both LLMs and SLMs lack of domain knowledge for trustworthy reasoning. Therefore, we propose ClinRaGen, enhancing SLMs by leveraging LLM-derived reasoning ability via rationale distillation and domain knowledge injection for trustworthy multimodal rationale generation. Key innovations include a sequential rationale distillation framework that equips SLMs with LLM-comparable mutlimodal reasoning abilities, and a knowledge-augmented attention mechanism that jointly unifies multimodal representation from time series and textual data in a same encoding space, enabling it naturally interpreted by SLMs while incorporating domain knowledge for reliable rationale generation. Experiments on real-world medical datasets show that ClinRaGen achieves state-of-the-art performance in disease diagnosis and rationale generation, demonstrating the effectiveness of combining LLM-driven reasoning with knowledge augmentation for improved interpretability.
Knowledge-Augmented Multimodal Clinical Rationale Generation for Disease Diagnosis with Small Language Models
Shuai Niu1, Jing Ma1, Hongzhan Lin1, Liang Bai2, Zhihua Wang3, Yida Xu1, Yunya Song1, and Xian Yang4thanks: * This is the corresponding author. Hong Kong Baptist University1, Shanxi University2 , Shanghai Institute for Advanced Study of Zhejiang University3, The University of Manchester4 {cssniu,majing,cshzlin}@comp.hkbu.edu.hk, zhihua.wang@zju.edu.cn, sxbailiang@126.com, {xuyida,yunyasong}@hkbu.edu.hk, xian.yang@manchester.ac.uk
@@ 1 Introduction
The widespread adoption of electronic health records (EHRs) has transformed deep learning applications in healthcare by providing diverse data modalities, including medical notes, laboratory (lab) test results, and clinical events. These multimodal inputs are crucial for disease diagnosis, mortality prediction, and drug discovery Niu et al. (2024); Laghuvarapu et al. (2024). Large language models (LLMs) have recently demonstrated strong diagnostic performance and reasoning capabilities through techniques such as prompt learning and Chain-of-Thought (CoT) reasoning Wei et al. (2022); Singhal et al. (2023); Chen et al. (2023). However, despite these advancements, LLMs face significant challenges in real-world clinical deployment due to high computational costs, the need for external domain-specific data integration, and difficulties in processing multimodal inputs—particularly numerical time-series lab test. More critically, LLMs lack the ability to generate clinically grounded multimodal rationales, limiting their interpretability in medical decision-making.
@@
To bring the best of both worlds, we propose ClinRaGen, a knowledge-augmented framework for multimodal clinical rationale generation. ClinRaGen enhances SLMs’ trustworthy mutlimodal reasoning capabilities from two aspects. First, it transfers LLM-derived reasoning to SLMs through a sequential rationale distillation paradigm. Second, unlike approaches that rely solely on LLM-generated rationales Kwon et al. (2024) or resource-intensive RAG Kang et al. (2024), we propose a knowledge-augmented attention mechanism that achieves dual functionality: Efficient integration of external medical knowledge to enable multimodal rationale generation grounded in clinical validity, ensuring the production of clinically meaningful explanations; Unification of time-series and textual EHRs within a shared encoding space, thereby enhancing multimodal representation learning and facilitates interpretable decision-making.
@@
  • We propose ClinRaGen, a multimodal framework that transfer LLM reasoning capabilities into SLMs for disease diagnosis and clinical rationale generation, achieving both accuracy and interpretability.
  • We introduce a knowledge-augmented attention mechanism that jointly encodes time-series EHRs into clinical textual representations while injecting domain knowledge, significantly improving mutlimodal rationale reliability and accuracy.
@@ 2 Related Work
Recent advancements in large-scale high-quality datasets and computational resources have enabled significant progress in Natural Language Processing (NLP), with improved training methodologies fueling the development of LLMs Touvron et al. (2023); Achiam et al. (2023). In healthcare, LLMs have been applied to clinical question answering and diagnostic reasoning Singhal et al. (2023); Yang et al. (2022). While effective in text-based tasks, these models struggle to generate clinically grounded multimodal rationales. Medical-specific LLMs Chen et al. (2023); Zhang et al. (2023) mitigate this issue through domain adaptation, but their high computational costs and reliance on large-scale training data limit scalability.
To improve efficiency, rationale distillation transfers LLM-derived reasoning ability to SLMs, reducing computational overhead while preserving interpretability Hsieh et al. (2023); Ho et al. (2023); Kang et al. (2024). Chain-of-thought prompting further enhances SLM reasoning capabilities Wei et al. (2022). However, most distillation approaches remain text-centric and lack robust multimodal EHRs integration Kang et al. (2024); Ho et al. (2023). RAG has been explored to improve rationale reliability by incorporating external knowledge, yet retrieval latency and adaptability remain key challenges Jiang et al. (2025). Despite these advancements, multimodal rationale generation remains an open challenge. Current models struggle to fuse textual, time-series, and structured medical knowledge into coherent clinical rationales.
@@ 3 Methodology
ClinRaGen consists of two key components: Knowledge Retrieval and LLM-Guided Rationale Generation (Section 3.1), which collects domain knowledge and generates LLM-derived rationales as distillation data for subsequent model training, and Knowledge-augmented Attention with Sequential Multimodal Rationale Distillation (Section 3.2), which progressively integrates structured knowledge to enhance multimodal reasoning in SLMs.
@@ 3.1.2 Generating Rationales for Distillation
Then, lab test-based rationales ($\bm{R}^{t}$) are generated using insights from both medical notes, time series anomalies, and the generate note-based rationales:
@@
Figure 4: Overview of ClinRaGen. (a) Model structure comprising a time series encoder, knowledge-augmented attention module, and a SLM. (b) Three-phase rationale distillation: Medical Note-based Rationale Distillation, Knowledge-Augmented Attention for Lab Test-based Rationale Distillation, and Multimodal Rationale Distillation.
@@ 3.2 Multimodal Rationale Distillation
Figure 4a presents the ClinRaGen framework, which comprises a Time Series Encoder for processing numerical lab test data, a Knowledge-Augmented Attention Module for integrating structured domain knowledge, and a SLM for generating disease diagnoses and structured clinical multimodal rationales. The framework enables progressive multimodal reasoning by leveraging structured knowledge and sequential learning mechanisms.
As illustrated in Figure 4b, ClinRaGen employs a three-phase rationale distillation paradigm that systematically integrates textual, numerical, and structured domain knowledge. The first phase distills medical note-based rationales, allowing the SLM to develop a foundational understanding of textual clinical information. The second phase introduces knowledge-augmented attention, aligning numerical lab test with structured medical knowledge to for distilling lab test-based rationales. The final phase fully integrates textual and numerical inputs, enabling the SLM to generate clinically coherent multimodal rationales to support disease diagnosis.
@@ 3.2.2 Phase 2: Knowledge Injection and Time-Series Rationale Distillation
To enable the SLM to effectively interpret numerical lab test data and generate time-series-based rationales ($\bm{R}^{t}$) that support disease diagnosis ($\bm{D}$), we introduce a Knowledge-Augmented Attention (KA) Module. This mechanism integrates domain-specific medical knowledge into the reasoning process, enhancing the model’s ability to produce clinically coherent and robust multimodal rationales.
@@
This phase ensures that the SLM can naturally interpret lab test while effectively leveraging medical knowledge to enhance its reasoning capabilities.
@@ 3.2.3 Phase 3: Full Multimodal Rationale Distillation
In the final phase, ClinRaGen is trained to generate full multimodal clinical rationales by integrating medical notes, lab test, and structured domain knowledge. To ensure effective multimodal reasoning, lab test $\bm{T}$ is formatted as prefix prompts Niu et al. (2024), allowing the SLM seamlessly incorporates it with textual EHRs.
@@ 4.1 Experimental Settings
Dataset: We evaluate ClinRaGen on two public EHR datasets: MIMIC-III Johnson et al. (2016) (28,456 EHRs include medical notes and time series lab tests) and MIMIC-IV Johnson et al. (2023) (28,900 EHRs). Both datasets use benchmark tools Harutyunyan et al. (2019) for time series processing, with missing values filled by nearest available data. We target 25 disease phenotypes and follow a 4:1 training-to-testing split Harutyunyan et al. (2019). Our model is available at github [^3] https://anonymous.4open.science/r/ClinRaGen-6C9D/.
Baseline Methods: To evaluate the effectiveness of ClinRaGen for disease diagnosis generation, we compared it with following baselines: Flan-T5 Chung et al. (2024), PROMPTEHR Wang and Sun (2022), FROZEN Tsimpoukelli et al. (2021), EHR-KnowGen Niu et al. (2024), Clinical CoT (with/without TSE) Kwon et al. (2024), and LLM-based models LLaMA-7B Touvron et al. (2023) (zero-shot and fine-tuning) and ChatGPT Open (2023) (zero-shot). Baseline and implementation details are provided in Appendices A.3 and A.4. For a fair comparison, all baselines (except LLaMA) use Flan-T5-Small as the backbone; our model employs Flan-T5-Small (ClinRaGen) and Flan-T5-Large (ClinRaGen*) for evaluate effect of varying scales. ChatGPT (GPT-3.5-turbo) serves as our teacher LLM. Results are averaged over five runs with statistical significance determined at p < 0.05 by t-test.
4.2 Disease Diagnosis Performance
Comparison with Baselines: We evaluate disease diagnosis using micro and macro precision, recall, and F1 scores. Table 1 shows that multimodal models outperform single-modality models, confirming the value of lab test results. Clinical CoT surpasses other baseline models, highlighting rationale distillation’s effectiveness. ClinRaGen (80M) achieves the best performance, with an average F1 score improvement of over 1.1% across all baselines, even outperforming LLaMA-7B-ft. Furthermore, ClinRaGen* (793M) improves by over 1.5%, significantly exceeding other baselines. The weak performance of zero-shot LLMs confirms the absence of data leakage. These results demonstrate ClinRaGen’s ability to match or surpass LLMs in clinical tasks through multimodal rationale distillation and the knowledge-augmented attention mechanism.
Figure 5: Model Parameter Counts and Micro F1 Scores
@@ 4.2 Disease Diagnosis Performance
Model Efficiency: We evaluate ClinRaGen’s efficiency by comparing model parameters, micro F1 scores (Figure 5), and training times (Table 3). ClinRaGen-Small achieves superior diagnostic performance with 80× fewer parameters and less than half the training time of LLaMA. These results highlight the effectiveness of our sequential multimodal distillation paradigm and KA mechanism in enabling efficient and accurate clinical reasoning.
4.3 Rationale Generation Performance
Evaluation Methods: To assess the quality of generated multimodal rationales and maximize the potential of SLMs, we evaluate ClinRaGen (80M) using five evaluation criteria—Correctness, Readability, Soundness, Consistency, and Persuasiveness—based on clinicians and prior research Lin et al. (2024); Kwon et al. (2024). Scores range from 1 to 5 on a Likert scale (details of criteria defined in Appendix A.5). We conduct both LLM-based and human evaluations. For LLM comparisons, we use Mistral-7B, LLaMA2-7B, and LLaMA3-8B with five-shot prompting. Distilled rationales from ChatGPT serve as ground truth (GT). Comparative LLMs receive time series anomalies and medical notes, while ClinRaGen directly processes numerical lab test and medical notes. Following Lin et al. (2024); Chiang and Lee (2023), we use GPT-4 to evaluate 1000 randomly selected samples. For human assessment, 15 professional postgraduates rate 100 samples, achieving moderate intra-class (0.637) and inter-class (0.608) agreement, indicating reasonable consistency despite the task’s subjectivity.
@@ 4.3 Rationale Generation Performance
Evaluation Results: Figures 6(a) and (b) show GPT-4 and human evaluations across five criteria, with closely aligned results. LLaMA3 performs best, benefiting from its large scale and pre-training. ClinRaGen ranks second, matching LLaMA3 in readability and correctness while surpassing LLaMA2 and Mistral, which often generate incoherent rationales. Unlike other LLMs relying on anomaly captions, ClinRaGen achieves the second-highest consistency score, demonstrating the KA mechanism’s effectiveness in consist multimodal reasoning. ClinRaGen also outperforms LLaMA2 and Mistral in soundness and persuasiveness, further underscoring our method’s effectiveness. Appendix A.6 further validates rationale quality using BLEU Papineni et al. (2002) and BERTScore Zhang et al. (2019).
Figure 6: Clinical rationale generation evaluation

### v5

@@ Abstract
Interpretation is critical for disease diagnosis, but existing models struggle to balance predictive accuracy with human-understandable rationales. While large language models (LLMs) offer strong reasoning abilities, their clinical use is limited by high computational costs and restricted multimodal reasoning ability. Small language models (SLMs) are efficient but lack advanced reasoning for integrating multimodal medical data. In addition, both LLMs and SLMs lack domain knowledge for trustworthy reasoning. Therefore, we propose ClinRaGen, enhancing SLMs by leveraging LLM-derived reasoning ability via rationale distillation and domain knowledge injection for trustworthy multimodal rationale generation. Key innovations include a sequential rationale distillation framework that equips SLMs with LLM-comparable multimodal reasoning abilities, and a knowledge-augmented attention mechanism that jointly unifies multimodal representation from time series and textual data in the same encoding space, enabling it to be naturally interpreted by SLMs while incorporating domain knowledge for reliable rationale generation. Experiments on real-world medical datasets show that ClinRaGen achieves state-of-the-art performance in disease diagnosis and rationale generation, demonstrating the effectiveness of combining LLM-driven reasoning with knowledge augmentation for improved interpretability.
@@ 1 Introduction
The widespread adoption of electronic health records (EHRs) has transformed deep learning applications in healthcare by providing diverse data modalities, including medical notes, laboratory (lab) test results, and clinical events. These multimodal inputs are crucial for disease diagnosis, mortality prediction, and drug discovery Niu et al. (2024b, a); Laghuvarapu et al. (2024). Large language models (LLMs) have recently demonstrated strong diagnostic performance and reasoning capabilities through techniques such as prompt learning and Chain-of-Thought (CoT) reasoning Wei et al. (2022); Singhal et al. (2023); Chen et al. (2023). However, despite these advancements, LLMs face significant challenges in real-world clinical deployment due to high computational costs, the need for external domain-specific data integration, and difficulties in processing multimodal inputs, particularly numerical time-series lab tests. More critically, LLMs lack the ability to generate clinically grounded multimodal rationales, limiting their interpretability in medical decision-making.
@@
To bring the best of both worlds, we propose ClinRaGen, a knowledge-augmented framework for multimodal clinical rationale generation. ClinRaGen enhances SLMs’ trustworthy multimodal reasoning capabilities from two aspects. First, it transfers LLM-derived reasoning to SLMs through a sequential rationale distillation paradigm. Second, unlike approaches that rely solely on LLM-generated rationales Kwon et al. (2024) or resource-intensive RAG Kang et al. (2024), we propose a knowledge-augmented attention mechanism that achieves dual functionality: Efficient integration of external medical knowledge to enable multimodal rationale generation grounded in clinical validity, ensuring the production of clinically meaningful explanations; Unification of time-series and textual EHRs within a shared encoding space, thereby enhancing multimodal representation learning and facilitates interpretable decision-making.
@@
  • We propose ClinRaGen, a multimodal framework that transfers LLM reasoning capabilities into SLMs for disease diagnosis and clinical rationale generation, achieving both accuracy and interpretability.
  • We introduce a knowledge-augmented attention mechanism that jointly encodes time-series EHRs into clinical textual representations while injecting domain knowledge, significantly improving multimodal rationale reliability and accuracy.
@@ 2 Related Work
Recent advancements in highly effective attention mechanisms Vaswani et al. (2017); Niu et al. (2021); Dao (2023); Yang et al. (2024), large-scale CoT datasets Cobbe et al. (2021); Zhong et al. (2023), improved computational resources have enabled significant progress in Natural Language Processing (NLP), with improved training methodologies fueling the development of LLMs Touvron et al. (2023); Achiam et al. (2023). In healthcare, LLMs have been applied to clinical question answering and diagnostic reasoning Singhal et al. (2023); Yang et al. (2022). While effective in text-based tasks, these models struggle to generate clinically grounded multimodal rationales. Medical-specific LLMs Chen et al. (2023); Zhang et al. (2023) mitigate this issue through domain adaptation, but their high computational costs and reliance on large-scale training data limit scalability.
To improve efficiency, rationale distillation transfers LLM-derived reasoning ability to SLMs, reducing computational overhead while preserving interpretability Lin et al. (2023); Hsieh et al. (2023); Ho et al. (2023); Kang et al. (2024). Chain-of-thought prompting further enhances SLM reasoning capabilities Wei et al. (2022). However, most distillation approaches remain text-centric and lack robust multimodal EHRs integration Kang et al. (2024); Ho et al. (2023). RAG has been explored to improve rationale reliability by incorporating external knowledge, yet retrieval latency and adaptability remain key challenges Jiang et al. (2025). Despite these advancements, multimodal rationale generation remains an open challenge. Current models struggle to fuse textual, time-series, and medical knowledge into coherent clinical rationales.
@@ 3 Methodology
ClinRaGen consists of two key components for the knowledge distillation: Knowledge Retrieval and LLM-Guided Rationale Generation (Section 3.1), which collects domain knowledge and generates LLM-derived rationales as distillation data for subsequent model training, and Multimodal Rationale Distillation (Section 3.2), which progressively integrates structured knowledge to enhance multimodal reasoning in SLMs.
@@ 3.1.2 Generating Rationales for Distillation
Then, lab test-based rationales ($\bm{R}^{t}$) are generated using insights from both medical notes, time series anomalies, and the generated note-based rationales:
@@
Figure 4: Overview of ClinRaGen. (a) Model structure comprising a time series encoder, a knowledge-augmented attention module, and an SLM. (b) Three-phase rationale distillation: Medical Note-based Rationale Distillation, Knowledge-augmented attention for Lab Test-based Rationale Distillation, and Multimodal Rationale Distillation.
@@ 3.2 Multimodal Rationale Distillation
Figure 4a presents the ClinRaGen framework, which comprises a Time Series Encoder for processing numerical lab test data, a knowledge-augmented attention module for integrating structured domain knowledge, and an SLM for generating disease diagnoses and structured clinical multimodal rationales. The framework enables progressive multimodal reasoning by leveraging structured knowledge and sequential learning mechanisms.
As illustrated in Figure 4b, ClinRaGen employs a three-phase rationale distillation paradigm that systematically integrates textual, numerical, and structured domain knowledge. The first phase distils medical note-based rationales, allowing the SLM to develop a foundational understanding of textual clinical information. The second phase introduces knowledge-augmented attention, aligning numerical lab tests with structured medical knowledge to distil lab test-based rationales. The final phase fully integrates textual and numerical inputs, enabling the SLM to generate clinically coherent multimodal rationales to support disease diagnosis.
@@ 3.2.2 Phase 2: Knowledge Injection and Time-Series Rationale Distillation
To enable the SLM to effectively interpret numerical lab test data and generate time-series-based rationales ($\bm{R}^{t}$) that support disease diagnosis ($\bm{D}$), we introduce a Knowledge-augmented Attention (KA) Module. This mechanism integrates domain-specific medical knowledge into the reasoning process, with the aim of enhancing the model’s focus, shifting it away from high-frequency but clinically irrelevant tokens towards clinically significant information. Consequently, this improved attention allows the model to produce more clinically coherent and robust multimodal rationales.
@@
This phase ensures that the SLM can naturally interpret lab tests while effectively leveraging medical knowledge to enhance its reasoning capabilities.
@@ 3.2.3 Phase 3: Full Multimodal Rationale Distillation
In the final phase, ClinRaGen is trained to generate full multimodal clinical rationales by integrating medical notes, lab tests, and structured domain knowledge. To ensure effective multimodal reasoning, lab test $\bm{T}$ is formatted as prefix prompts Niu et al. (2024a), allowing the SLM to seamlessly incorporate it with textual EHRs.
@@ 4.1 Experimental Settings
Dataset: We evaluate ClinRaGen on two public EHR datasets: MIMIC-III Johnson et al. (2016) (28,456 EHRs include medical notes and time series lab tests) and MIMIC-IV Johnson et al. (2023) (28,900 EHRs). Both datasets use benchmark tools Harutyunyan et al. (2019) for time series processing, with missing values filled by the nearest available data. We target 25 disease phenotypes and follow a 4:1 training-to-testing split Harutyunyan et al. (2019). Our model is available at github [^3] https://github.com/Healthcare-Data-Mining-Laboratory/ClinRaGen.
Baseline Methods: To evaluate the effectiveness of ClinRaGen for disease diagnosis generation, we compared it with following baselines: Flan-T5 Chung et al. (2024), PROMPTEHR Wang and Sun (2022), FROZEN Tsimpoukelli et al. (2021), EHR-KnowGen Niu et al. (2024a), Clinical CoT (with/without TSE) Kwon et al. (2024), and LLM-based models LLaMA-7B Touvron et al. (2023) (zero-shot and fine-tuning) and ChatGPT Open (2023) (zero-shot). Baseline and implementation details are provided in Appendices A.3 and A.4. For a fair comparison, all baselines (except LLaMA) use Flan-T5-Small as the backbone; our model employs Flan-T5-Small (ClinRaGen) and Flan-T5-Large (ClinRaGen*) to evaluate the effect of varying scales. ChatGPT (GPT-3.5-turbo) serves as our teacher LLM. Results are averaged over five runs, with statistical significance determined at p < 0.05 by t-test.
@@
Figure 5: Model Parameter Counts and Micro F1 Scores
4.2 Disease Diagnosis Performance
Comparison with Baselines: We evaluate disease diagnosis using micro and macro precision, recall, and F1 scores. Table 1 shows that multimodal models outperform single-modality models, confirming the value of lab test results. Clinical CoT surpasses other baseline models, highlighting the rationale for distillation’s effectiveness. ClinRaGen (80M) achieves the best performance, with an average F1 score improvement of over 1.1% across all baselines, even outperforming LLaMA-7B-ft. Furthermore, ClinRaGen* (793M) improves by over 1.5%, significantly exceeding other baselines. The weak performance of zero-shot LLMs confirms the absence of data leakage. These results demonstrate ClinRaGen’s ability to match or surpass LLMs in clinical tasks through multimodal rationale distillation and the knowledge-augmented attention mechanism. Appendix A.6 provides further experimental details, revealing a substantial improvement of over 3% in F1 scores achieved by our method across different SLMs and through distillation from varied LLMs.
Model Efficiency: We evaluate ClinRaGen’s efficiency by comparing model parameters, micro F1 scores (Figure 5), and training times (Table 3). ClinRaGen (80M) achieves superior diagnostic performance with 80× fewer parameters and less than half the training time of LLaMA-7B. These results highlight the effectiveness of our sequential multimodal distillation paradigm and Knowledge-augmented attention mechanism in enabling efficient and accurate clinical reasoning.
@@ 4.2 Disease Diagnosis Performance
4.3 Rationale Evaluation with LLM-as-a-Judge and Human
Evaluation Methods: To assess the quality of generated multimodal rationales and maximize the potential of SLMs, we evaluate ClinRaGen (80M) using five evaluation criteria: Correctness, Readability, Soundness, Consistency, and Persuasiveness, as informed by clinicians and prior research Lin et al. (2024); Kwon et al. (2024). These criteria are scored on a Likert scale ranging from 1 to 5 (details of the criteria are provided in Appendix A.5). We conduct both LLM-based evaluations (using LLM-as-a-Judge Gu et al. (2024)) and human evaluations. For LLM comparisons, we use Mistral-7B, LLaMA2-7B, and LLaMA3-8B with five-shot prompting. Distilled rationales from ChatGPT serve as ground truth (GT). Comparative LLMs receive time series anomalies and medical notes, while ClinRaGen directly processes numerical lab test and medical notes. Following Lin et al. (2024); Chiang and Lee (2023), we use GPT-4 to evaluate 1000 randomly selected samples. For human assessment, 15 professional postgraduates rate 100 samples, achieving moderate intra-class (0.637) and inter-class (0.608) agreement, indicating reasonable consistency despite the task’s subjectivity.
Evaluation Results: Figures 6(a) and (b) show GPT-4 and human evaluations across five criteria, with closely aligned results. LLaMA3 performs best, benefiting from its large scale and pre-training. ClinRaGen ranks second, matching LLaMA3 in readability and correctness while surpassing LLaMA2 and Mistral, which often generate incoherent rationales. Unlike other LLMs relying on anomaly captions, ClinRaGen achieves the second-highest consistency score, demonstrating the knowledge-augmented attention mechanism’s effectiveness in consistent multimodal reasoning. ClinRaGen also outperforms LLaMA2 and Mistral in soundness and persuasiveness, further underscoring our method’s effectiveness.
4.4 Rationale Evaluation with BLUE and BERTScore.
Figure 6: Rationale evaluation with LLMs and Human